# From AI for Science to Agentic Science: A Survey on Autonomous Scientific Discovery

### v1

@@ 8.1 General Frameworks and Methodologies
In astronautics, the LLMSat agent was developed to serve as a high-level, goal-oriented controller for autonomous spacecraft, aiming to reduce reliance on human mission control for deep space exploration [182]. Tested in the Kerbal Space Program simulator, the work found that while current LLMs have limitations in handling high-complexity missions, their performance can be improved with advanced prompting frameworks and by carefully defining the agent’s level of authority over the spacecraft. Finally, in a more theoretical application, the CosmoAgent system uses LLM-based agents to simulate interactions between human and hypothetical extraterrestrial civilizations [298]. By programming agents with different worldviews and ethical paradigms, the system explores potential inter-civilizational dynamics, providing a novel tool for studying cooperation and conflict under conditions of asymmetric information.

### v2

@@ 8.1 General Frameworks and Methodologies
In astronautics, the LLMSat agent was developed to serve as a high-level, goal-oriented controller for autonomous spacecraft, aiming to reduce reliance on human mission control for deep space exploration [233]. Tested in the Kerbal Space Program simulator, the work found that while current LLMs have limitations in handling high-complexity missions, their performance can be improved with advanced prompting frameworks and by carefully defining the agent’s level of authority over the spacecraft. In fundamental physics, the ArgoLOOM framework was developed to act as an agentic AI orchestrator that autonomously coordinates computational tools across cosmology, collider, and nuclear physics domains [20]. Tested on case studies involving sterile‑neutrino scenarios, the system demonstrated its ability to link large‑scale cosmological simulations with collider and deep‑inelastic‑scattering analyses through an LLM‑driven planning pipeline. In experimental physics, the Accelerator Assistant framework was developed as an agentic AI system capable of autonomously executing multi‑stage experiments at a large‑scale synchrotron user facility [121]. Deployed at the Advanced Light Source, it translates natural‑language prompts into structured execution plans that integrate data retrieval, control‑system interaction, and analysis workflows under strict safety and reproducibility constraints, demonstrating a two‑order‑of‑magnitude reduction in experiment preparation time for machine‑physics studies. Finally, in a more theoretical application, the CosmoAgent system uses LLM-based agents to simulate interactions between human and hypothetical extraterrestrial civilizations [380]. By programming agents with different worldviews and ethical paradigms, the system explores potential inter-civilizational dynamics, providing a novel tool for studying cooperation and conflict under conditions of asymmetric information.